OpenStudy (anonymous):

Find the derivative of f(x) = 2 divided by x at x = -2.

OpenStudy (anonymous):

\[f(x)=\frac{ 2}{ x }=2x^{-1}\] Do you know the power rule?

OpenStudy (anonymous):

wouldn't it be -1?

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

no, what did you get for f'(x)?

OpenStudy (anonymous):

ohhhh 1/2

OpenStudy (anonymous):

2/(-2)^2

OpenStudy (anonymous):

yes 1/2

OpenStudy (anonymous):

What about f(x)= -9/x, x=-8 =9/64 @peachpi

OpenStudy (anonymous):

wait actually 64/9 since nine is negative

OpenStudy (anonymous):

you were right 9/64

OpenStudy (anonymous):